OpenStudy (anonymous):

The area of a sector of a circle with a central angle of 2 rad is 25 m2. Find the radius of the circle.

OpenStudy (p0sitr0n):

do proportions: 25/2=x/2pi. solve for x to get the area. then use area=pi*r^2 to solve for r